OpenStudy (anonymous):

compute the following 325C1

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

How many different four-digit numbers can be made using the digits 1, 2, 3, 4, 5, 6 if no digit can be used more than once?

OpenStudy (wolf1728):

I think the answer to the combination question is 6*5*4*3*2*1 which I think is 720

OpenStudy (wolf1728):

MY MISTAKE 6*5*4*3

OpenStudy (anonymous):

COMPUTE THE FOLLOWING; 8c8

OpenStudy (wolf1728):

and c means combination?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

CN U HELP ?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (wolf1728):

Looking at wikipedia to find exactly how to do that :-)

OpenStudy (anonymous):

ty

jimthompson5910 (jim_thompson5910):

In general n C n = 1 where n is any integer

jimthompson5910 (jim_thompson5910):

*any positive integer

jimthompson5910 (jim_thompson5910):

this is because there is only one way to have a group of n people (drawn from a total pool of n people) and order doesn't matter

jimthompson5910 (jim_thompson5910):

so what this means is that you replace n with 8 to get 8 C 8 = 1

OpenStudy (anonymous):

10c4 * 4c2 = ? @jim_thompson5910

OpenStudy (wolf1728):

or basically it is asking how many k items can be chosen from a group of n

OpenStudy (wolf1728):

10c4 means how many ways can k things (4) be selected from n things (10). combinations = n! / (n-k)! * k! 10c4 = 10*9*8*7*6*5*4*3*2*1 / 6*5*4*3*2*1 * (4*3*2*1) = 10*9*8*7 / 4*3*2*1 = 10*3*7 =210 4c2 = 4! / 2! * 2! 4c2 = 3*2*1 = 6 so, 10c4 * 4c2 = 210*6 = 1,260

OpenStudy (anonymous):

TY so much

OpenStudy (wolf1728):

u r welcome